# Global Convergence of Policy Gradient for Entropy Regularized Linear-Quadratic Control with multiplicative noise

### v1

@@ Appendix B Proofs in Section 3
$f(K,\Sigma)-f(K^{*},\Sigma^{*})\geq\frac{\mathbb{E}[x_{0}^{2}]}{S_{K}^{2}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\nabla f^{\top}(K,\Sigma)\nabla f(K,\Sigma)$ (45)
@@ C.3 Two useful Lemmas
For the second term, $\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)$. We want to invoke the Vector Bernstein Inequality to show that with high probability $\hat{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)||<\frac{\epsilon}{2}$. Consider the sample $i^{th}$ of a single path $K+U_{i}$ and observe that $\|\frac{n}{r_{1}^{2}}f(K+U_{i})U_{i}\|\leq\frac{2nf(K,\Sigma)}{r_{1}}$ if we assume $f(K+U_{i},\Sigma)\leq 2f(K,\Sigma)$. Also,
@@
$\widetilde{\nabla}_{\Sigma}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r^{2}}f(K,\Sigma+V_{i})V_{i}$
is $\epsilon$ close to $\nabla_{\Sigma}f(K,\Sigma)$ with high probability under some conditions. We break $\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)$ into two terms,
$\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)$ $=(\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)-\nabla_{\Sigma}f(K,\Sigma))+(\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma))$ (58)
For the first term, by Lemma $4.8$ we have $\|\nabla_{\Sigma}f(K,\Sigma+V)-\nabla_{\Sigma}f(K,\Sigma)\|<\frac{\epsilon}{2}$ if $r_{2}\leq\frac{\epsilon}{2h_{9}}$. And because $\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)=\mathbb{E}[\nabla_{\Sigma}f(K,\Sigma+V)]$, we have $\|\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)-\nabla_{\Sigma}f(K,\Sigma+V)\|_{F}\leq\frac{\epsilon}{2}$ by triangle inequality. For the second term $\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)$. Consider a single sample $V_{j}$ from the distribution. Assume $f(K,\Sigma+V_{j})\leq 2f(K,\Sigma)$, then $\|\frac{n}{r_{2}^{2}}f(K,\Sigma+V_{j})V_{j}\|\leq\frac{2nf(K,\Sigma)}{r_{2}}$. Also, note that $\mathbb{E}\left[\frac{n}{r_{2}^{2}}f(K,\Sigma+V_{j})V_{j}\right]=\mathbb{E}\left[\widetilde{\nabla}_{\Sigma}\right]=\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)$. Now, given that $\mathbb{E}\left[\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\right]=0$, $\|\frac{n}{r_{2}^{2}}f(K,\Sigma+V_{j})V_{j}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\|_{F}\leq\frac{2nf(K,\Sigma)}{r_{2}}+\frac{\epsilon}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$, and if $\mathbb{E}[\|\frac{n}{r^{2}}f(K,\Sigma+V_{j})V_{j}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\|_{F}^{2}]=\sigma_{V}^{2}\leq\left(\frac{2nf(K,\Sigma)}{r_{2}}\right)^{2}+\left(\frac{\epsilon}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}\right)^{2}$ then by the Vector Bernstein Inequality we have $\mathbb{P}\big(\big\|\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\big\|\geq\frac{\epsilon}{2}\big)\leq n\cdot\exp{\big(-\frac{3}{2}\cdot\frac{\frac{\epsilon^{2}}{4}M}{3\sigma_{V}^{2}+R_{2}\frac{\epsilon}{2}}\big)}\leq\kappa_{2}$, $R_{2}=\frac{2nf(K,\Sigma)}{r_{2}}+\frac{\epsilon}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$, and $\kappa_{2}\in(0,1)$. This also gives us that the minimum samples: $M\geq\log\left(\frac{d}{\kappa_{2}}\right)\cdot\frac{4}{\epsilon^{2}}\left(2\sigma_{V}^{2}+R_{2}\frac{\epsilon}{3}\right)$.
@@
$\nabla_{\Sigma}^{\prime}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r^{2}}f^{(l)}(K,\Sigma+V_{j})V_{j}.$
We break $\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)$ into three terms,
$\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)=(\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}^{\prime})+(\nabla_{\Sigma}^{\prime}-\hat{\nabla}_{\Sigma})+(\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma))$
For the first term, we have
$\sum_{j=0}^{l-1}Q(x_{j}^{i})^{2}+(u_{j}^{i})^{\top}R(u_{j}^{i})+\tau\log\pi(u_{j}|x_{j})$ $\leq\Gamma\mathbb{E}\left[\sum_{j=0}^{l-1}Qx_{j}^{2}+u_{j}^{\top}Ru_{j}+\tau\log\pi(u_{j}|x_{j})\right]$
$\leq\Gamma\mathbb{E}\left[\sum_{j=0}^{\infty}Qx_{j}^{2}+u_{j}^{\top}Ru_{j}+\tau\log\pi(u_{j}|x_{j})\right]$
$\leq\Gamma L^{2}f(K,\Sigma+V_{i})$
$\leq 2\Gamma L^{2}f(K,\Sigma)$
Where $|x_{0}^{i}|\leq L$. Then since $\|\nabla_{\Sigma}^{\prime}\|\leq\frac{\epsilon}{3}+\|\widehat{\nabla}_{\Sigma}\|\leq\frac{5\epsilon}{6}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$ so the norm of each sample is bounded, $\mathbb{E}[\widehat{\nabla}_{\Sigma}-\nabla^{\prime}_{\Sigma}]$ = 0, and assuming the variance is bounded as well then by the Vector Bernstein Inequality $\mathbb{P}\left(\|\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}^{\prime}\|>\frac{\epsilon}{3}\right)\leq d\cdot\exp{\left(-\frac{3}{2}\cdot\frac{\frac{\epsilon^{2}}{9}M}{3\sigma_{V^{\prime}}^{2}+R_{2}^{\prime}\frac{\epsilon}{3}}\right)}\leq\kappa_{2}$ if $M\geq\frac{6}{\epsilon^{2}}\log\left(\frac{d}{\kappa_{2}}\right)\left(3\sigma_{V^{\prime}}^{2}+R_{2}^{\prime}\frac{\epsilon}{3}\right)$, so $\|\widehat{\nabla}_{\Sigma}-\nabla^{\prime}_{\Sigma}\|\leq\frac{\epsilon}{3}$ with probability at least $1-\kappa_{2}$. For the second term, we have
[…]
$\leq\frac{1}{M}\cdot\frac{n}{r_{2}}\sum_{i=1}^{M}\gamma^{l}\Bigg|\left(|Q|+\|K\|^{2}\|R\|\right)\frac{f(K,\Sigma+V_{i})-\left(Tr((\Sigma+V_{i})R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma+V_{i}|)\right)}{|Q|}\Bigg|$
$+\Bigg|\frac{Tr((\Sigma+V_{i})R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma+V_{i}|)}{1-\gamma}\Bigg|$
$\leq\frac{1}{M}\cdot\frac{n}{r_{2}}\sum_{i=1}^{M}\gamma^{l}\left[\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}\right)2f(K,\Sigma)\right$.
$\quad+\left.\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\right)\left|Tr((\Sigma+V_{i})R)-\frac{\tau}{2}\left(n+2\log(2\pi)^{n}|\Sigma|\right)\right|\right]$
$\leq\frac{1}{M}\cdot\frac{n}{r_{2}}M\gamma^{l}\big[\big(1+\frac{\|K\|^{2}\|R\|}{|Q|}\big)2f(K,\Sigma)$
$\quad+\big(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\big)\left(\left|2Tr(\Sigma R)\right|+\frac{\tau}{2}\left(n+2\log(2\pi)^{n}|\Sigma|\right)\right)\big]$
$\leq\frac{\epsilon}{3},$
if $|Tr(\Sigma+V_{i})|\leq 2|Tr(\Sigma)|$, $\|\Sigma+V_{i}\|\leq 2\|\Sigma\|$, and $l\geq$
$\frac{\log\left(\frac{\epsilon r_{2}}{3n}\right){\log(\gamma)}-\log\left[\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}\right)2f(K,\Sigma)+\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\right)(|\text{Tr}(\Sigma R)|+\frac{\tau}{2}|(n+2\log(2\pi)^{n}|\Sigma|)|)\right]}{\log(\gamma)}$
@@
For the third term, based on the previous proof, we have $\|\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with high probability $1-\kappa_{2}$ under conditions given in the lemma. ∎

### v3

@@ Appendix B Proofs in Section 3
$f(K,\Sigma)-f(K^{*},\Sigma^{*})\geq\frac{\mathbb{E}[x_{0}^{2}]}{S_{K}^{2}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\nabla f^{\top}(K,\Sigma)\nabla f(K,\Sigma).$
@@ C.3 Two useful Lemmas
For the second term, $\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)$. We want to invoke the Vector Bernstein Inequality to show that with high probability $\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)||<\frac{\epsilon}{2}$. Consider the sample $i^{th}$ of a single path $K+U_{i}$ and observe that $\|\frac{n}{r_{1}^{2}}f(K+U_{i})U_{i}\|\leq\frac{2nf(K,\Sigma)}{r_{1}}$ if we assume $f(K+U_{i},\Sigma)\leq 2f(K,\Sigma)$. Also,
@@
Define:
$\widehat{\nabla}_{L}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]V_{i}$
@@
$\widetilde{\nabla}_{L}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}$
$\|\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with high probability under some conditions. We break $\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)$ into two terms,
$\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)$ $=(\nabla_{L}f_{r_{2}}(K,L)-\nabla_{L}f(K,\Sigma))+(\widetilde{\nabla}_{L}-\nabla_{L}f_{r_{2}}(K,L))$ (27)
For the first term, by Lemma $4.5$ we have $\|\nabla_{L}f(K,(L+V)(L+V)^{\top})-\nabla_{L}f(K,\Sigma)\|<\frac{\epsilon}{6}$ if $r_{2}\leq\frac{\epsilon}{6h_{9}}$. And because $\nabla_{L}f_{r_{2}}(K,L)=\mathbb{E}[\nabla_{L}f(K,(L+V)(L+V)^{\top})]$, we have
$\|\nabla_{L}f_{r_{2}}(K,L)-\nabla_{L}f(K,\Sigma)\|_{F}\leq\frac{\epsilon}{6}$
by triangle inequality.
For the second term $\widetilde{\nabla}_{L}-\nabla_{L}f_{r_{2}}(K,L)$. Note that $\mathbb{E}\left[\widetilde{\nabla}_{L}\right]=\sum_{i=1}^{M}\|\frac{2n(n+1)}{2r_{2}^{2}}f(K,\Sigma)V_{i}\|=\nabla_{L}f_{r_{2}}(K,L)$, using $f(K,\Sigma_{i})\leq 2f(K,\Sigma)$ we have
$\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}\|\leq\|\frac{2n(n+1)}{2r_{2}^{2}}f(K,\Sigma)V_{i}\|\leq\frac{n(n+1)}{r_{2}}f(K,\Sigma)$
using the the result of the first term and triangle inequality we have
$\|\nabla_{L}f_{r_{2}}(K,L)\|_{F}\leq\|\nabla_{L}f(K,\Sigma)\|_{F}+\frac{\epsilon}{6}$
$\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}-\nabla_{L}f_{r_{2}}(K,L)\|_{F}\leq\frac{n(n+1)}{r_{2}}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{\epsilon}{6}$
so we have
$\mathbb{E}[\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}-\nabla_{L}f_{r_{2}}(K,\Sigma)\|_{F}^{2}]\leq R_{Z_{1}}:=\frac{n(n+1)}{r_{2}}f(K,\Sigma)+\overline{\|\nabla_{L}f_{r_{2}}(K,L)\|}+\frac{\epsilon}{6}.$
Next we consider
$\mathbb{E}((\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i})^{\top}\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}))-\nabla_{L}f_{r_{2}}(K,L)^{\top}\nabla_{L}f_{r_{2}}(K,L)$
$\quad\leq\max_{V_{i}}\|\frac{n(n+1)}{2r_{2}^{2}}f(K,\Sigma_{i})V_{i}\|_{F}-\|\nabla_{L}f_{r_{2}}(K,L)\|_{F}^{2}$
$\quad\leq\sigma_{Z_{1}}:=\frac{n(n+1)}{r_{2}}f(K,\Sigma)$
Using Vector Bernstein Inequality, we know that if
$M\geq\frac{2}{(\epsilon/3)^{2}}(\sigma_{Z_{1}}^{2}+\frac{R_{Z_{1}}{\epsilon/3}}{3})\log(\frac{1}{1-\sqrt{1-\kappa}})$
then $\|\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with probability at least $1-\kappa$.
@@
$\nabla_{L}^{\prime}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n(n+1)}{2r_{2}^{2}}f^{(l)}(K,\Sigma_{i})V_{i}.$
We break $\widehat{\nabla}_{L}-\nabla_{L}f(K,\Sigma)$ into three terms,
$\widehat{\nabla}_{L}-\nabla_{L}f(K,\Sigma)=(\widehat{\nabla}_{L}-\nabla_{L}^{\prime})+(\nabla_{L}^{\prime}-\widetilde{\nabla}_{L})+(\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma))$
@@
For the third term, based on the previous proof, we have $\|\widetilde{\nabla}_{L}-\nabla_{L}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with probability $1-\kappa_{2}$ under conditions given in the lemma. Furthermore, we have $\|\widetilde{\nabla}_{L}\|\leq\|\nabla_{L}f(K,\Sigma)\|+\frac{\epsilon}{3}$
For the second term, we have
$\|\nabla_{L}^{\prime}-\widetilde{\nabla}_{L}\|$ $=\frac{n(n+1)}{2r_{2}^{2}}\frac{1}{M}\sum_{i=1}^{M}\left(f^{(l)}(K,\Sigma_{i})-f(K,\Sigma_{i})\right)\|V_{i}\|$
$\leq\frac{n(n+1)}{2r_{2}}\frac{1}{M}\sum_{i=1}^{M}\left(f^{(l)}(K,\Sigma_{i})-f(K,\Sigma_{i})\right)$
If set
$l\geq\frac{\log\frac{2r_{2}\epsilon}{3n(n+1)}-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{log\gamma}$
then $\|\nabla_{L}^{\prime}-\widetilde{\nabla}_{L}\|\leq\frac{\epsilon}{3}$ by lemma 4.2. Furthermore, by triangle inequality we have
$\|\nabla_{L}^{\prime}\|\leq\|\widetilde{\nabla}_{L}\|+\frac{\epsilon}{3}\leq\|\nabla_{L}f(K,\Sigma)\|+\frac{2\epsilon}{3}\leq\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{2\epsilon}{3}$
For the first term, we will use Vector Bernstein inequality as $\mathbb{E}[\widehat{\nabla}_{L}]=\nabla^{\prime}_{L}$. By the assumption we have
$\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}|x_{t})$ $\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$
$\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{\infty}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$
$\leq\Gamma L^{2}f(K,\Sigma_{i})$
$\leq 2\Gamma L^{2}f(K,\Sigma),$
where $|x_{0}^{i}|\leq L$. Define $\widehat{\nabla}^{i}_{L}$ so that $\widehat{\nabla}_{L}=\sum_{i=1}^{M}\widehat{\nabla}^{i}_{L}$
$\|\widehat{\nabla}^{i}_{L}-{\nabla}_{L}^{\prime}\|_{F}$ $\leq\|\widehat{\nabla}^{i}_{L}\|_{F}+\|{\nabla}_{L}^{\prime}\|_{F}$
$\leq\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)+\|{\nabla}_{L}^{\prime}\|_{F}$
$\leq R_{Z}:=\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)+\overline{\|\nabla_{L}f(K,\Sigma)\|}+\frac{2\epsilon}{3}$
$\mathbb{E}[(\widehat{\nabla}^{i}_{L})^{\top}\widehat{\nabla}^{i}_{L}]-{\nabla_{L}^{\prime}}^{\top}\nabla_{L}^{\prime}$
$\quad\leq\max_{V_{i}}\|\widehat{\nabla}^{i}_{L}\|^{2}_{F}-\|\nabla_{L}^{\prime}\|_{F}^{2}$
$\quad\leq\|\frac{n(n+1)}{2r_{2}^{2}}2\Gamma L^{2}f(K,\Sigma)V_{i}\|^{2}-\|\nabla_{L}^{\prime}\|_{F}^{2}$
$\quad\leq\big(\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)\big)^{2}-\|\nabla_{L}^{\prime}\|_{F}^{2}$
$\quad\leq\sigma_{Z_{2}}:=\big(\frac{n(n+1)}{r_{2}}\Gamma L^{2}f(K,\Sigma)\big)^{2}$
If
$M>\frac{2n(n+1)}{2(\epsilon/3)^{2}}\big(\sigma_{{Z_{2}}}^{2}+\frac{R_{{Z_{2}}}\epsilon}{9\sqrt{n(n+1)/2}}\big)\log\frac{n(n+1)/2+1}{1-\sqrt{1-\kappa}}$
then
$\|\widehat{\nabla}_{L}-\nabla^{\prime}_{L}\|<\frac{\epsilon}{3}$
with probability $1-\sqrt{1-\kappa}$. Combining the above we have
$\|\widehat{\nabla}_{L}-\nabla_{L}f(K,\Sigma)\|\leq\epsilon$
with probability at least $1-\kappa$. As
$\nabla_{L}f(K,\Sigma=LL^{\top})$ $=\nabla_{L}(LL^{\top})\nabla_{\Sigma}f(K,\Sigma)$
$=2L\nabla_{\Sigma}f(K,\Sigma)$
Now take $\epsilon=\frac{\epsilon_{2}}{2\|L^{-1}\|}$, we have
$\|\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)\|=\|(2L)^{-1}\widehat{\nabla}_{\Sigma}-(2L)^{-1}\nabla_{\Sigma}f(K,\Sigma)\|\leq\epsilon_{2}$
with probability $1-\kappa$